A copper wire is 2.50 m long and has a diameter of 2.5 mm. How much current will it draw when connected to 3.5 V source?


1
Expert's answer
2022-04-22T13:53:29-0400
"I=\frac{V}{R}=\frac{V}{\rho\frac{l}{A}}"

"I=\frac{3.5}{1.68*10^{-8}*\frac{2.5}{3.14*0.0025^2/4}}=409\:\rm A"